OpenStudy (anonymous):

What is the pH of a 0.256 M NaOH, please explain how you get the answer

OpenStudy (anonymous):

0.1 M HCl has [H+] = 0.10 M, [OH-] = 1x 10-13 M] because ,[H+] [OH-] = 1x 10-14 so, 0.25 M NaOH has [OH-] = 0.25 = 25 x 10-2 . as [H+] [OH-] = 1x 10-14 => [H+} = 1x 10-14/ [OH-] => [H+] = 1x 10-14/25 x 10-2 = 4.0x10-14 M pH = - log(H+) = -log(4.0x10-14 )=> pH =13.40, pOH = 14-13.40 = 0.6